OpenStudy (anonymous):

The period of a mechanical wave is 12 seconds. What is the frequency of the wave?

OpenStudy (ajprincess):

Frequency=1/period Does that help?

OpenStudy (konradzuse):

T = 1/f T is the period.

OpenStudy (anonymous):

so if the period is 12 the frequency is 12?

OpenStudy (konradzuse):

no

OpenStudy (konradzuse):

f *T = 1/f *f (f * T)/T = 1/T f = 1/T

OpenStudy (anonymous):

if T is the period (12 seconds) and its = to \[\frac{ 1 }{ frequency }\] it would look like \[12=\frac{ 1 }{ f }\] right?

OpenStudy (konradzuse):

yes

OpenStudy (anonymous):

12f=1 ; 1/12 ; \[\frac{ 1 }{ 12 } \] or 0.083

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (konradzuse):

mhm